OpenStudy (anonymous):

given that f(x) = 2x + 1 and g(x) = x^2 -2.find (f o g) (x + 2). choices a. 2x^2 - 8x + 5 c. 2x^2 + 8x - 5 b. 2x^2 + 8x +5 d.2x^2 - 8x - 5 pls help for my H.w

OpenStudy (anonymous):

discrete math functions

OpenStudy (anonymous):

I know this as composite func, f o g means f is function of g, so plug g into x in f

OpenStudy (anonymous):

your correct composite function

OpenStudy (anonymous):

what is the correct sign negative or positive?

OpenStudy (anonymous):

\[f(x)=2x+1\] \[ g(x) = x^2 -2\] \[(f o g)=2(x^2-2)+1\]